In a certain federal prison, it is known that 2/3 of the inmates are under 25 years of age. It is also known that 3/5 of the inmates are male and that 5/8 of the inmates are female or 25 years of age or older. What is the probability that a prisoner selected at random from this prison is female and at least 25 years old?


1
Expert's answer
2021-06-16T17:19:16-0400

Let "A" denotes the event "inmate is under 25 years of age".

Let "M" denotes the event "inmate is a male".

Let "F" denotes the event "inmate is a female".

Given "P(A)=\dfrac{2}{3}, P(M)=\dfrac{3}{5}, P(F\cup A')=\dfrac{5}{8}."


"P(A')=1-P(A)=1-\dfrac{2}{3}=\dfrac{1}{3}"

"P(F)=1-P(M)=1-\dfrac{3}{5}=\dfrac{2}{5}"


"P(F\cap A')=P(F)+P(A')-P(F\cup A')"

"P(F\cap A')=\dfrac{2}{5}+\dfrac{1}{3}-\dfrac{5}{8}=\dfrac{48+40-75}{120}"

"=\dfrac{13}{120}"